OpenStudy (anonymous):

A jet plane lands with a speed of 112 m/s and can accelerate at a maximum rate of −5.40 m/s2 as it comes to rest. (a) From the instant the plane touches the runway, what is the minimum time needed before it can come to rest? (b) Can this plane land on a small tropical island airport where the runway is 0.800 km long?

Parth (parthkohli):

The three equations of motion with uniform acceleration \(a\):\[v_f = v_i + at\]\[d = v_i t + \frac{1}{2}at^2\]\[v_f^2 = v_i^2 + 2ad\]

Parth (parthkohli):

For the first question, you need an equation of motion with the following variables: 1. Initial velocity (= 112 m/s) 2. Final velocity (= 0) 3. Acceleration (= -5.4 m/s2) 4. Time (unknown) Which one would it be? :)

OpenStudy (anonymous):

first one....?

Parth (parthkohli):

That's correct. What variables can you use for the second question?

OpenStudy (anonymous):

distance

Parth (parthkohli):

yes, and what variables do you have known to you?

OpenStudy (anonymous):

I got it!!! Thank you! 20.74 s and No... Your equations helped a lot :)